# Exact Hull Reformulation for Quadratically Constrained Generalized Disjunctive Programs

### v1

@@ 6.4 Constrained layout problem
$\min_{\delta_{x},\delta_{y},\mathbf{x},\mathbf{y},\mathbf{W},\mathbf{Y}}\quad$ $\sum_{i,j\in N,\,i<j}c_{ij}\cdot\|(\delta_{x_{ij}},\;\delta_{y_{ij}})\|_{p}$
s.t. $\delta_{x_{ij}}\geq x_{i}-x_{j}\quad$ $i,j\in N,\,i<j$
$\delta_{x_{ij}}\geq x_{j}-x_{i}\quad$ $i,j\in N,\,i<j$
$\delta_{y_{ij}}\geq y_{i}-y_{j}\quad$ $i,j\in N,\,i<j$
$\delta_{y_{ij}}\geq y_{j}-y_{i}\quad$ $i,j\in N,\,i<j$
$\left[\begin{array}[]{c}Y_{ij}^{1}\\ x_{i}+L_{i}/2\leq x_{j}-L_{j}/2\end{array}\right]\vee\left[\begin{array}[]{c}Y_{ij}^{2}\\ x_{j}+L_{j}/2\leq x_{i}-L_{i}/2\end{array}\right]$ (20)
$\quad\vee\left[\begin{array}[]{c}Y_{ij}^{3}\\ y_{i}+H_{i}/2\leq y_{j}-H_{j}/2\end{array}\right]\vee\left[\begin{array}[]{c}Y_{ij}^{4}\\ y_{j}+H_{j}/2\leq y_{i}-H_{i}/2\end{array}\right]$ $i,j\in N,\,i<j$ (25)
$\bigvee_{t\in T}\left[\begin{array}[]{c}W_{it}\\ (x_{i}+L_{i}/2-x_{ct})^{2}+(y_{i}+H_{i}/2-y_{ct})^{2}\leq r_{t}^{2}\\ (x_{i}+L_{i}/2-x_{ct})^{2}+(y_{i}-H_{i}/2-y_{ct})^{2}\leq r_{t}^{2}\\ (x_{i}-L_{i}/2-x_{ct})^{2}+(y_{i}+H_{i}/2-y_{ct})^{2}\leq r_{t}^{2}\\ (x_{i}-L_{i}/2-x_{ct})^{2}+(y_{i}-H_{i}/2-y_{ct})^{2}\leq r_{t}^{2}\end{array}\right]$ $i\in N$ (31)
$Y_{ij}^{1}\underline{\vee}Y_{ij}^{2}\underline{\vee}Y_{ij}^{3}\underline{\vee}Y_{ij}^{4}\quad$ $i,j\in N,\,i<j$
$\underset{t\in T}{\underline{\bigvee}}W_{it},\quad$ $i\in N$
$x_{i}^{l}\leq x_{i}\leq x_{i}^{u}\quad$ $i\in N$
$y_{i}^{l}\leq y_{i}\leq y_{i}^{u}\quad$ $i\in N$
$\delta_{x_{ij}},\delta_{y_{ij}}\in\mathbb{R}_{+}\quad$ $i,j\in N,\,i<j$
$x_{i},y_{i}\in\mathbb{R}\quad$ $i\in N$
$Y_{ij}^{1},Y_{ij}^{2},Y_{ij}^{3},Y_{ij}^{4}\in\{\text{False},\text{True}\}\quad$ $i,j\in N,\,i<j$
$W_{it}\in\{\text{False},\text{True}\}\quad$ $i\in N,\,t\in T$
@@
To assess reformulation performance, we tested six instances of this problem implemented in the Pyomo.GDP example library [^3] https://github.com/Pyomo/pyomo/tree/main/examples/gdp/constrained_layout using both $\ell_{1}$ and $\ell_{2}$ norm. A summary of results for two representative instances (CLay0305 using $\ell_{2}$ norm and CLay0304 using $\ell_{1}$ norm) is provided in Table 3.
For CLay0305, the proposed exact Hull Reformulation solution time was comparable to or greater compared to $\varepsilon$-approximation; however, in the case of SCIP, the $\varepsilon$-approximation reported an incorrect result with large objective and lower bound values. This is attributed to numerical instabilities introduced by the approximation, consistent with observations from other problem types.
In contrast, for CLay0304, the exact Hull Reformulation demonstrated faster solve times with BARON and SCIP compared to the $\varepsilon$-approximation, while in the case of Gurobi, although the solution time was similar, the $\varepsilon$-approximation was faster. With BARON, the $\varepsilon$-approximation hit the 1800s limit with an 11% gap remaining, whereas the exact Hull Reformulation solved the problem in 11.39s to optimality.
Results for all six instances using both $\ell_{1}$ and $\ell_{2}$ norms are provided in Appendix D. While the limited size and number of benchmark problems prevent drawing general conclusions, the observed results suggest that, overall, the exact Hull Reformulation performs on par with the $\varepsilon$-approximation for constrained layout problems.
Table 3: Solver outcomes for CLay0305 and CLay0304 under different norms
(a) CLay0305 with $\ell_{2}$ norm
Strategy Time (s) Objective Lower Bound
@@
BigM 53.57 6594.21 6594.21
Hull ($\varepsilon$-approx.) 61.94 6594.21 6594.21
Hull Exact 72.13 6594.21 6594.21
Binary Mult. 78.77 6594.21 6594.21
Solver: Gurobi
BigM 42.85 6594.21 6594.21
Hull ($\varepsilon$-approx.) 218.42 6594.21 6594.21
Hull Exact 858.55 6594.22 6594.21
Binary Mult. 78.53 6594.21 6594.21
@@
BigM 30.50 6594.21 6594.21
Hull ($\varepsilon$-approx.) 21.79 41171.96 41171.96
Hull Exact 20.73 6594.21 6594.21
Binary Mult. 28.77 6594.21 6594.21
(b) CLay0304 with $\ell_{1}$ norm
Strategy Time (s) Objective Lower Bound
Solver: BARON
BigM 0.70 40262.39 40262.39
Hull ($\varepsilon$-approx.) 1808.30 40790.97 36280.70
Hull Exact 11.39 40262.39 40262.39
Binary Mult. 3.84 40262.39 40262.39
Solver: Gurobi
BigM 0.79 40262.39 40262.39
Hull ($\varepsilon$-approx.) 1.79 40262.39 40262.36
Hull Exact 2.13 40262.37 40262.36
Binary Mult. 0.78 40262.39 40262.39
Solver: SCIP
BigM 1.41 40262.39 40262.39
Hull ($\varepsilon$-approx.) 9.59 40262.39 40262.39
Hull Exact 3.28 40262.39 40262.39
Binary Mult. 2.84 40262.39 40262.39
@@ 7 Conclusions
This paper proposed a strategy to form an exact Hull Reformulation for any quadratic constraints in GDP problems, effectively addressing the drawbacks associated with the traditional $\varepsilon$-approximation of the closure of the perspective function method. By preserving the quadratic functional structure of constraints and providing an exact reformulation without approximations, the proposed approach improves solver performance, numerical stability, and reliability.
Theoretical results establish that the proposed reformulation is equivalent to the original exact Hull Reformulation, yielding an optimization problem with an identical feasible region. While the conventional exact Hull Reformulation is not expressed in a closed analytical form, our approach provides an analytically defined reformulation that is exact for quadratic functions and therefore directly implementable in practice.
Extensive computational experiments demonstrated that the proposed exact Hull Reformulation approach outperformed the $\varepsilon$-approximation approach in various problem types, including random GDP instances, CSTR network optimization, $k$-means clustering, and constrained layout problems. The results highlighted not only a generally improved solution time, but also a reduction in numerical instabilities and solver failures to provide the correct result.
Furthermore, it is expected that the results obtained could be further improved if the convexity detection methods of the solvers could recognize certain reformulations of the quadratically constrained GDPs as convex, or if the solvers could directly utilize the convexity information obtained before the reformulation. This suggests an area for future development to improve the solvers’ performance.
Several future research directions emerge from this work. First, the performance of the extension of the exact Hull Reformulation to polynomial constraints of higher degree could be thoroughly evaluated, as we did for quadratic constraints in this work. Second, a systematic evaluation could be conducted to compare the direct application of Hull Reformulation to non-convex GDPs against strategies that first convexify the constraints. Third, it would be valuable to compare the performance of the Hull Reformulation with alternative reformulation methods, such as Big-M or Binary Multiplication, when these are applied directly to non-convex GDPs.
Overall, the findings of this study strongly support the adoption of the exact Hull Reformulation as the standard approach for GDP optimization problems involving quadratic constraints.
@@ Acknowledgments
The authors acknowledge and express sincere gratitude to the Davidson School of Chemical Engineering at Purdue University for providing a supportive research environment and funding. The authors thank the GAMS Development Corporation for providing a personal license and access to the BARON solver for our group. We thank Gurobi Optimization for providing access to software through their academic licensing program. We also thank Prof. Can Li for providing access to the hardware.
Appendices
@@ Appendix A From Binary Multiplication to the Hull Reformulation
We outline how the Hull Reformulation for a single disjunction can be systematically derived from the simpler Binary Multiplication reformulation by taking a convex combination of disjuncts. This derivation does not rely on convexity assumptions, demonstrating that the Hull Reformulation is applicable to both convex and non-convex GDPs. It also offers additional insight into the interpretation of the reformulation when applied to non-convex problems. This appendix is adapted from Appendix A of [31].
@@
Binary Multiplication reformulation
@@
Taking a convex combination
Let $\mathbf{x}_{i}$ be a copy of the original variable associated with the feasible region of disjunct $i$. We now form a convex combination of the disjuncts, which yields the convex hull of the disjunction if all $h_{i}(\mathbf{x})$ are convex. Note, however, that convex combinations can also be taken for non-convex sets, but the resulting set will not, in general, be convex, which is illustrated in Figure 1. Thus, these steps remain valid for non-convex problems, although they no longer guarantee the convex hull of the disjunction. The convex combination of disjuncts can then be defined as follows:
@@
Collecting all constraints yields the Hull Reformulation of the disjunction:
@@
$\left(\text{cl}\,\widetilde{h}_{i}\right)(\mathbf{v}_{i},y_{i})\leq 0,,i\in D$
@@
Thus, Hull Reformulation can be systematically obtained from Binary Multiplication by taking a convex combination of disjuncts in lifted space. For non-convex $h_{i}(\mathbf{x})$, the construction remains valid, but the continuous relaxation is no longer guaranteed to be convex.
@@ Appendix B Extension of the Exact Hull Reformulation to Polynomial Constraints of Degree $d$
The continuous relaxation of the standard Hull Reformulation for this polynomial constraint introduces the following set in the lifted space:
@@
which, in the case of polynomial constraints, simplifies to:
@@
Therefore, when applying Hull Reformulation, any polynomial disjunctive constraint $\sum_{k=0}^{d}p_{k}(\mathbf{x})\leq 0$, can be reformulated in a way that forms exact hull and preserves polynomial structure as $\sum_{k=0}^{d}p_{k}(\mathbf{v})\,y^{\,d-k}\leq 0$ (sum of monomials of degree $d$), instead of using the closure of the perspective function or its $\varepsilon$‐approximation.
@@
Appendix C Continuously Stirred Tank Reactor (CSTR) Network
This appendix provides a description and the set of equations used to formulate the Continuously Stirred Tank Reactor (CSTR) network optimization problem considered in Section 6, adapted from prior formulations in the literature [28, 9].
Each stage $n\in N=\{1,\dots,N_{T}\}$ may operate as a CSTR or be bypassed; exactly one stage receives the fresh feed and exactly one stage receives the recycle stream returned from the outlet of the last stage. Component indices are collected in $I=\{A,B\}$. Variables include molar flow rates $F_{i,n}$ and $FR_{i,n}$, total flow rates $Q_{n}$ and $Q_{FR,n}$, reactor volumes $V_{n}$ (with volume cost $c_{n}$), reaction rates $r_{i,n}$, the recycle flow $(R_{i},Q_{R})$, the product flow $(P_{i},Q_{P})$, and an auxiliary variables $t_{n}$ are used to reduce cubic terms to quadratic form. Boolean variables $YF_{n}$, $YR_{n}$, and $YP_{n}$ denote, respectively, that fresh feed enters stage $n$, recycle enters stage $n$, and stage $n$ operates as a CSTR. All continuous variables are non-negative.
Complete GDP formulation can be presented as follows:
$\min_{F_{i,n},FR,r,t,Q,V,c,YF_{n},YR_{n},YP_{n}}$ $\sum_{n\in N}c_{n}$ (32)
s.t. $F_{i,NT}-F0_{i}-FR_{i,NT}-r_{i,NT}V_{NT}=0$ $\forall\ i\in I$
$Q_{NT}-Q_{F0}-Q_{FR,NT}=0$
$F_{i,n}-F_{i,n+1}-FR_{i,n}-r_{i,n}V_{n}=0$ $\forall\ n\in N\setminus\{N_{T}\},\;i\in I$
$Q_{n}-Q_{n+1}-Q_{FR,n}=0$ $\forall\ n\in N\setminus\{N_{T}\}$
$F_{i,1}-P_{i}-R_{i}=0$ $\forall\ i\in I$
$Q_{1}-Q_{P}-Q_{R}=0$
$P_{i}Q_{1}-F_{i,1}Q_{P}=0$ $\forall\ i\in I$
$0\displaystyle 95\,Q_{P}-P_{B}=$0
$V_{n}-V_{n-1}=0$ $\forall\ n\in N\setminus\{1\}$
$Q_{n}^{2}-t_{n}=0$ $\forall\ n\in N$
$\left[\begin{aligned} &YP_{n}\\ &r_{A,n}t+kF_{A,n}F_{B,n}=0\\ &r_{B,n}+r_{A,n}=0\\ &c_{n}-V_{n}=0\end{aligned}\right]\operatorname*{\underline{\bigvee}}\left[\begin{aligned} &\neg YP_{n}\\ &FR_{i,n}=0\quad(i\in I)\\ &r_{i,n}=0\quad(i\in I)\\ &Q_{FR,n}=0\\ &c_{n}=0\end{aligned}\right]$ $\forall\ n\in N$
$\left[\begin{aligned} &YR_{n}\\ &FR_{i,n}-R_{i}=0\quad(i\in I)\\ &Q_{FR,n}-Q_{R}=0\end{aligned}\right]\operatorname*{\underline{\bigvee}}\left[\begin{aligned} &\neg YR_{n}\\ &FR_{i,n}=0\quad(i\in I)\\ &Q_{FR,n}=0\end{aligned}\right]$ $\forall\ n\in N$
$\underline{\bigvee}_{n\in N}YF_{n},\qquad\underline{\bigvee}_{n\in N}YR_{n}$
$YP_{n}\iff\left(\bigwedge_{j\in\{1,2,..n\}}\neg YF_{j}\right)\vee YF_{n}$ $\forall\ n\in N$
$YR_{n}\Longrightarrow YP_{n}$ $\forall\ n\in N$
$F_{i,n},FR_{i,n},P_{i},R_{i},Q_{n},Q_{FR,n},Q_{F0}\geq 0$ $\forall\ n\in N,i\in I$
$Q_{P},Q_{R},V_{n},c_{n},r_{i,n},t\geq 0$ $\forall\ n\in N,i\in I$
$YF_{n},YR_{n},YP_{n}\in\{\text{False},\text{True}\}$ $\forall\ n\in N$
Appendix D Constrained Layout Problem Results
This appendix presents detailed results for the six instances of the constrained layout problem used in the computational experiments discussed in Section 6. These instances originate from the Pyomo.GDP example library and include variations in both objective norms ($\ell_{1}$ and $\ell_{2}$) and layout configurations.
Table 4 summarizes solver outcomes, specifically, the number of optimal solutions, timeouts, and incorrect results, for each of the four reformulation strategies (Big-M, Hull with $\varepsilon$-approximation, exact Hull Reformulation, and Binary Multiplication) across three solvers (BARON, Gurobi, and SCIP). These results provide an overview of the performance and reliability of each reformulation method.
Figures 10 and 11 provide visual comparisons of performance. Figure 10 presents the number of instances solved up to a given wall‑clock. Figure 11 directly compares solution times for each problem instance between the proposed exact Hull Reformulation and the $\varepsilon$-approximation approach.
The proposed exact Hull Reformulation generally matched or outperformed the $\varepsilon$-approximation in terms of solution correctness and numerical stability. In two cases, the $\varepsilon$-approximation produced incorrect solutions, when using SCIP, and in one instance Gurobi reported a wrong solution in the case of the proposed exact Hull Reformulation.
BARON and SCIP showed similar performance in terms of solution time for $\varepsilon$-approximation and exact Hull Reformulation, while in the case of Gurobi $\varepsilon$-approximation showed better results.
Overall, although the limited number and size of the benchmark instances preclude broad generalizations, these results demonstrate that the proposed exact Hull Reformulation is generally competitive with the $\varepsilon$-approximation, and the results may vary depending on the solver.

### v2

@@ 4.4 Constrained layout problem (convex GDP)
$\min_{\delta_{x},\delta_{y},\mathbf{x},\mathbf{y},\mathbf{W},\mathbf{Y}}$ $\sum_{i,j\in N,\,i<j}c_{ij}\,\|(\delta_{x_{ij}},\delta_{y_{ij}})\|_{p}$ (24)
s.t. $\delta_{x_{ij}}\geq x_{i}-x_{j},\qquad\forall\,i,j\in N,\ i<j,$
$\delta_{x_{ij}}\geq x_{j}-x_{i},\qquad\forall\,i,j\in N,\ i<j,$
$\delta_{y_{ij}}\geq y_{i}-y_{j},\qquad\forall\,i,j\in N,\ i<j,$
$\delta_{y_{ij}}\geq y_{j}-y_{i},\qquad\forall\,i,j\in N,\ i<j,$
$\left[\begin{array}[]{c}Y_{ij}^{1}\\ x_{i}+L_{i}/2\leq x_{j}-L_{j}/2\end{array}\right]\vee\left[\begin{array}[]{c}Y_{ij}^{2}\\ x_{j}+L_{j}/2\leq x_{i}-L_{i}/2\end{array}\right]$
$\vee\left[\begin{array}[]{c}Y_{ij}^{3}\\ y_{i}+H_{i}/2\leq y_{j}-H_{j}/2\end{array}\right]\vee\left[\begin{array}[]{c}Y_{ij}^{4}\\ y_{j}+H_{j}/2\leq y_{i}-H_{i}/2\end{array}\right],$
$\hskip 20.00003pt\forall\,i,j\in N,\ i<j,$
$\bigvee_{t\in T}\left[\begin{array}[]{c}W_{it}\\ (x_{i}+L_{i}/2-x_{ct})^{2}+(y_{i}+H_{i}/2-y_{ct})^{2}\leq r_{t}^{2}\\ (x_{i}+L_{i}/2-x_{ct})^{2}+(y_{i}-H_{i}/2-y_{ct})^{2}\leq r_{t}^{2}\\ (x_{i}-L_{i}/2-x_{ct})^{2}+(y_{i}+H_{i}/2-y_{ct})^{2}\leq r_{t}^{2}\\ (x_{i}-L_{i}/2-x_{ct})^{2}+(y_{i}-H_{i}/2-y_{ct})^{2}\leq r_{t}^{2}\end{array}\right],$
$\hskip 20.00003pt\forall\,i\in N,$
$Y_{ij}^{1}\underline{\vee}Y_{ij}^{2}\underline{\vee}Y_{ij}^{3}\underline{\vee}Y_{ij}^{4},\qquad\forall\,i,j\in N,\ i<j,$
$\underset{t\in T}{\underline{\bigvee}}W_{it},\qquad\forall\,i\in N,$
$x_{i}^{l}\leq x_{i}\leq x_{i}^{u},\qquad\forall\,i\in N,$
$y_{i}^{l}\leq y_{i}\leq y_{i}^{u},\qquad\forall\,i\in N,$
$\delta_{x_{ij}},\delta_{y_{ij}}\in\mathbb{R}_{+},\qquad\forall\,i,j\in N,\ i<j,$
$x_{i},y_{i}\in\mathbb{R},\qquad\forall\,i\in N,$
$Y_{ij}^{1},Y_{ij}^{2},Y_{ij}^{3},Y_{ij}^{4}\in\{\mathrm{False},\mathrm{True}\},\qquad\forall\,i,j\in N,\ i<j,$
$W_{it}\in\{\mathrm{False},\mathrm{True}\},\qquad\forall\,i\in N,\ t\in T.$
@@
To assess reformulation performance, we tested six instances of this problem implemented in the Pyomo.GDP example library [^3] https://github.com/Pyomo/pyomo/tree/main/examples/gdp/constrained_layout using both $\ell_{1}$ and $\ell_{2}$ norm. A summary of results is provided in Figures 7 and Table 4.
While the limited size and number of benchmark problems prevent drawing general conclusions, the observed results suggest that, overall, both CEHR and GEHR performed on par with the $\varepsilon$-approximation for constrained layout problems.
(a) Gurobi
(b) BARON
(c) SCIP
Strategy Optimal Timeout Objective mismatch Total
Solver: Gurobi
BigM 12 0 0 12
Hull $\varepsilon$-approx. ($\varepsilon=10^{-4}$) 10 0 2 12
GEHR 10 0 2 12
CEHR 10 2 0 12
Binary Mult. 11 1 0 12
@@
BigM 12 0 0 12
Hull $\varepsilon$-approx. ($\varepsilon=10^{-4}$) 11 0 1 12
GEHR 12 0 0 12
CEHR 10 2 0 12
Binary Mult. 12 0 0 12
@@
BigM 12 0 0 12
Hull $\varepsilon$-approx. ($\varepsilon=10^{-4}$) 12 0 0 12
GEHR 12 0 0 12
CEHR 12 0 0 12
Binary Mult. 12 0 0 12
@@ 5 Conclusions
This paper investigated HR for GDPs with quadratic disjunctive constraints, with the goal of avoiding the convexity recognition, numerical and relaxation-quality drawbacks introduced by the standard $\varepsilon$-approximation of the closure of the perspective function.
Hull Reformulation for non-convex GDPs.
HR is a valid reformulation technique even when a GDP contains non-convex constraints. The reformulation preserves the intended on/off behavior at binary values, recovering the original constraints when $y=1$ and rendering them redundant when $y=0$.
In the non-convex setting, however, the continuous relaxation produced by HR is not guaranteed to be convex. Moreover, while HR is designed to represent the convex hull of the feasible region of each individual disjunction when the constraints in the disjuncts are convex, this interpretation does not hold when the disjunctive constraints are non-convex. Consequently, the resulting relaxation should not be interpreted as producing the convex hull of the disjunction.
As a result, HR of non-convex GDP is naturally targeted to global MINLP solvers, which may additionally construct convex underestimators internally when computing bounds.
General Exact Hull Reformulation for quadratic constraints.
For general quadratic constraints, possibly non-convex, we showed GEHR that preserves quadratic degree and avoids the fractional expressions introduced by the $\varepsilon$-approximation. Also, because the reformulation is exact, it avoids feasible-region enlargement associated with $\varepsilon>0$ and removes the need to tune an approximation parameter.
Conic Exact Hull Reformulation for convex quadratics.
For convex quadratic disjunctive constraints ($Q\succeq 0$), we showed derivation of CEHR, which is expressed by a rotated second-order cone representable inequality of the form $\mathbf{v}^{\top}Q\mathbf{v}\leq t\,y$, together with a linear inequality linking $(t,\mathbf{v},y)$. This provides a solver-friendly reformulation for convex quadratics with a conic-representable structure, enabling direct exploitation by conic-capable solvers.
Summary of computational findings.
Across random convex and non-convex quadratic GDPs, convex $k$-means clustering, a non-convex CSTR network benchmark, and constrained layout instances, the proposed exact formulations reduced numerical failures relative to the $\varepsilon$-approximation and frequently improved or had comparable runtime.
In our experiments, CEHR was consistently the most reliable hull formulation on convex benchmarks and achieved the best overall performance, especially for solvers that recognize and exploit conic structure. Its substantial advantage over the $\varepsilon$-approximation is largely attributable to the fact that the $\varepsilon$-approximation can obscure convexity and hinder convexity detection by the solver.
On both convex and non-convex benchmarks, GEHR was more numerically stable and typically faster than the $\varepsilon$-approximation unless it was a convex GDP and a solver was explicitly informed about convexity.
Implications for modeling.
Overall, our results support using CEHR for convex quadratically constrained GDPs and GEHR for non-convex instances over $\varepsilon$-approximation when HR is applied to GDP.
Future work.
This work motivates several directions for future research. First, the polynomial generalization in Appendix B should be benchmarked systematically to understand when preserving polynomial degree yields practical gains over alternative lifting strategies. Second, for non-convex GDPs, it would be valuable to compare direct application of hull reformulations with approaches that convexify disjunctive constraints prior to reformulation, and to benchmark hull reformulations against other formulations, such as Big-M, when applied directly to non-convex GDPs. Finally, improving solver-side convexity detection and handling of perspective-type (i.e., $\varepsilon$-approximation) constraints remains an important opportunity for future solver development.
@@ Acknowledgments
The authors acknowledge and express sincere gratitude to the Davidson School of Chemical Engineering at Purdue University for providing a supportive research environment and funding.
The authors thank the GAMS Development Corporation for providing a personal license and access to the BARON solver for our group.
The authors thank Gurobi Optimization for providing access to software through their academic licensing program.
The authors thank Prof. Can Li for providing access to the hardware.
The authors thank Kevin Furman, Nicolas Sawaya, Ignacio Grossmann, Nikolaos Sahinidis for their helpful feedback provided through discussions of the preprint and poster presentations.
This material is based upon work supported by the National Science Foundation under Award No. 2430617.
@@ Appendix A From Binary Multiplication to the Hull Reformulation
We outline how HR for a single disjunction can be systematically derived from the simpler Binary Multiplication reformulation by taking a convex combination of disjuncts. This derivation does not rely on convexity assumptions, demonstrating that HR is applicable to both convex and non-convex GDPs. It also offers additional insight into the interpretation of the reformulation when applied to non-convex problems. This appendix is adapted from Appendix A of Lee and Grossmann [19].
@@
A.1 Binary Multiplication reformulation
@@
A.2 Taking a convex combination
Let $\mathbf{x}_{i}$ be a copy of the original variable associated with the feasible region of disjunct $i$. We now form a convex combination of the disjuncts, which yields the convex hull of the disjunction if all $h_{i}(\mathbf{x})$ are convex. Note, however, that we don’t have to assume convexity and convex combinations can also be taken for non-convex sets, but the resulting set will not, in general, be convex, which is illustrated in Figure 1. Thus, these steps remain valid for non-convex problems, although they no longer guarantee the convex hull of the disjunction. The convex combination of disjuncts can then be defined as follows:
@@ A.3 Hull Reformulation
Collecting all constraints yields HR of the disjunction:
@@
$\left(\text{cl}\,\widetilde{h}_{i}\right)(\mathbf{v}_{i},y_{i})\leq 0,i\in D$
@@
Thus, HR can be systematically obtained from Binary Multiplication by taking a convex combination of disjuncts in lifted space. Note that convexity is not assumed at any step, and for non-convex function $h_{i}(\mathbf{x})$, the construction remains valid, but the continuous relaxation is no longer guaranteed to be convex.
@@ Appendix B Extension of the Exact Hull Reformulation to Polynomial Constraints of Degree $d$
The continuous relaxation of the standard HR for this polynomial constraint introduces the following set in the lifted space:
@@
which, in the case of polynomial constraints simplifies to:
@@
Therefore, when applying HR, any polynomial disjunctive constraint $\sum_{k=0}^{d}p_{k}(\mathbf{x})\leq 0$, can be reformulated in a way that forms exact hull and preserves polynomial structure as $\sum_{k=0}^{d}p_{k}(\mathbf{v})\,y^{\,d-k}\leq 0$ (sum of monomials of degree $d$), instead of using $\varepsilon$‐approximation.
@@
Appendix C Algebraic second-order cone representations for the Conic Exact Hull
Section 3.2 shows that HR of convex quadratic disjunctive constraints admits an exact second-order cone representation (CEHR). In Pyomo.GDP, constraints are passed to solvers in algebraic form, so conic-capable solvers can exploit the structure only if they recognize it from the provided algebraic representation. Since solver interfaces and presolve routines may prefer different but equivalent representations of the same cone constraint, we tested several equivalent forms of the Conic Exact Hull on the random convex benchmark set.
Derivation of the conic perspective constraint.
Consider a convex quadratic inequality in a disjunct:
[…]
In HR, the closure of the perspective function constraint (for $y>0$) is
$\frac{\mathbf{v}^{\mathsf{T}}Q\,\mathbf{v}}{y}+\mathbf{c}^{\mathsf{T}}\mathbf{v}+d\,y\;\leq\;0,$ (26)
together with the standard disaggregation bounds $\mathbf{x}^{\ell}y\leq\mathbf{v}\leq\mathbf{x}^{u}y$ (and $y\in[0,1]$). Introduce an epigraph variable $t$ to isolate the quadratic-over-$y$ term:
$\mathbf{v}^{\mathsf{T}}Q\,\mathbf{v}\;\leq\;t\,y,\qquad t+\mathbf{c}^{\mathsf{T}}\mathbf{v}+d\,y\;\leq\;0,\qquad t\geq 0,\;y\geq 0.$ (27)
For $y>0$, (27) is equivalent to (26) by dividing the first inequality by $y$ and substituting into the second. At $y=0$, the disaggregation bounds force $\mathbf{v}=\mathbf{0}$, so $\mathbf{v}^{\mathsf{T}}Q\mathbf{v}=0$ and (27) implies $t=0$, matching the closure value $\bigl(\operatorname{cl}\widetilde{h}\bigr)(\mathbf{0},0)=0$.
Factorized form.
Since $Q\succeq 0$, there exists a matrix $L$ such that
$Q\;=\;L\,L^{\mathsf{T}}.$ (28)
Therefore,
$\mathbf{v}^{\mathsf{T}}Q\,\mathbf{v}\;=\;\|L^{\mathsf{T}}\mathbf{v}\|_{2}^{2},$ (29)
and the rotated-cone part of (27) can be written as $\|L^{\mathsf{T}}\mathbf{v}\|_{2}^{2}\leq t\,y$ with $t\geq 0$ and $y\geq 0$.
Tested reformulation representations (names match the plot legend).
All representations below impose the same linear constraint $t+\mathbf{c}^{\mathsf{T}}\mathbf{v}+d\,y\leq 0$, the same nonnegativity conditions $t\geq 0$, $y\geq 0$, and the same disaggregation bounds $\mathbf{x}^{\ell}y\leq\mathbf{v}\leq\mathbf{x}^{u}y$. They differ only in the algebraic form used for the rotated second-order cone component.
  • Conic Exact Hull (No factorization). Direct quadratic-over-product inequality:
    $\mathbf{v}^{\mathsf{T}}Q\,\mathbf{v}\;\leq\;t\,y.$ (30)
  • Conic Exact Hull (Only factorized). Using (28)–(29):
    $\|L^{\mathsf{T}}\mathbf{v}\|_{2}^{2}\;\leq\;t\,y.$ (31)
  • Conic Exact Hull (No Sqrt, No Extra Var). Starting from (31), use the identity $(t+y)^{2}-(t-y)^{2}=4ty$ to obtain an equivalent quadratic inequality (for $t\geq 0$, $y\geq 0$):
    $4\|L^{\mathsf{T}}\mathbf{v}\|_{2}^{2}+(t-y)^{2}\;\leq\;(t+y)^{2},\qquad t\geq 0,\;y\geq 0.$ (32)
  • Conic Exact Hull (Sqrt, No Extra Var). Taking square roots of (32) yields the standard SOC form:
     $\sqrt{\,4\|L^{\mathsf{T}}\mathbf{v}\|_{2}^{2}+(t-y)^{2}\,}\;\leq\;t+y,\qquad t\geq 0,\;y\geq 0.$ (33)
  • Conic Exact Hull (Sqrt, Extra Var). Introduce auxiliary variables
    $\mathbf{z}=2L^{\mathsf{T}}\mathbf{v},\qquad w=t-y,\qquad s=t+y,\qquad s\geq 0,$ (34)
    and impose the SOC
    $\sqrt{\|\mathbf{z}\|_{2}^{2}+w^{2}}\;\leq\;s.$ (35)
  • Conic Exact Hull (No Sqrt, Extra Var). Using the same auxiliaries (34), impose the squared SOC form
    $\|\mathbf{z}\|_{2}^{2}+w^{2}\;\leq\;s^{2},\qquad s\geq 0.$ (36)
Explicit rotated-cone membership (optional).
For reference, (31) is equivalent to rotated second-order cone membership:
$\left(\frac{t}{2},\,y,\,L^{\mathsf{T}}\mathbf{v}\right)\in\mathcal{Q}_{r}\;:=\;\left\{(\alpha,\beta,\mathbf{w})\ \middle|\ 2\alpha\beta\geq\|\mathbf{w}\|_{2}^{2},\ \alpha\geq 0,\ \beta\geq 0\right\}.$
Computational comparison and choice of default representation.
Figure 8 compares these Conic Exact Hull representations on the random convex benchmark set using Gurobi, BARON, and SCIP. The figure also includes BigM as a baseline reference.
For Gurobi and SCIP, which have explicit conic capabilities, Conic Exact Hull (No factorization), denoted as just Conic Exact Hull in Figure 8, performs among the best representations, suggesting that both solvers recognize and exploit the rotated-cone structure directly from this algebraic form.
For BARON, Conic Exact Hull (No factorization) performs comparably to most other Conic Exact Hull representations on this benchmark. One representation performs noticeably better here, namely Conic Exact Hull (Sqrt, No Extra Var), but the same representation performs poorly for Gurobi and SCIP.
Taken together, these results support using Conic Exact Hull (No factorization) in the direct form (30) as the overall default representation: it is consistently competitive and is the strongest option for solvers that actively exploit conic structure (Gurobi and SCIP), while remaining comparable to alternatives for BARON.